# Integrating Multi-Physics Simulations and Machine Learning to Define the Spatter Mechanism and Process Window in Laser Powder Bed Fusion

### v1

@@ 2.1 Simulation of LPBF via OpenFOAM
where m is the mass transfer between phases, $C^{\prime}$ is Lee’s model constant, $T$ is the current temperature, $T_{A}$ is the solidification or melting temperature, L is the latent heat. It is worth noting that the same set of equation works for both solidification ($T$ $<$ $T_{A}$ and $C^{\prime}$ negative) and melting ($T$ $>$ $T_{A}$ and $C^{\prime}$ positive) [45].
Table 1: Material Parameters used to simulate the SS316L melting process in OpenFOAM simulation with subscript $s$ for solid phase, $l$ for liquid phase, $v$ for vapor and $n$ for shielding gas nitrogen
@@
2.2 Machine Learning
2.2.1 Data Collection
The spatter and the meltpool features were extracted using ParaView, an open-source, multi-platform scientific data analysis and visualization tool. The extracted features were used to validate our developed tracking algorithm which drastically accelerate the process. Detecting trace of variations in the property distributions between the spatter and the meltpool can enhance classification task. We implement different ML models as discussed below to identify the high performing model for the spatter/meltpool classification. Investigating the dataset for classification task will aid the understanding of spatter formation mechanism.
2.2.2 The Extra Trees Classifier (Extremely Randomized Trees)
The Extra Trees Classifier (Extremely Randomized Trees) is an ensemble learning method fundamentally similar to Random Forests [46]. However, it differs in the way it constructs the trees. In Extra Trees, randomness goes one step further in the way splits are computed. The top-down splitting in the tree learner is randomized. Instead of computing the most discriminative thresholds, thresholds are drawn at random for each candidate feature, and the best of these randomly generated thresholds is picked as the splitting rule. This often allows the model to reduce variance, at the cost of a slight increase in bias.
2.2.3 Extreme Gradient Boosting classifier (XGBClassifier)
EXtreme Gradient Boosting classifier (XGBClassifier) is an implementation of gradient boosted decision trees designed for speed and performance [47]. It is known for its efficiency at handling large datasets and its effectiveness in improving model performance through boosting. XGBoost also includes a unique feature called regularization which helps in reducing overfitting.
2.2.4 Light Gradient Boosting Machine (LGBMClassifier)
Light Gradient Boosting Machine (LGBMClassifier) is a gradient boosting framework that uses tree-based learning algorithms. It is designed for distributed and efficient training, particularly on large datasets [48]. LGBM is known for its high efficiency, handling large data, and lower memory usage. Its efficiency comes from histogram-based optimizations in the tree learning process.
2.2.5 Random Forest Classifier (RFClassifier)
Random Forest Classifier is an ensemble learning method that operates by constructing a multitude of decision trees at training time and outputting the class that is the mode of the classes of the individual trees [49]. Random forests correct for decision trees’ habit of overfitting to their training set.
2.2.6 BaggingClassifier
BaggingClassifier is a meta-estimator ensemble learning method that fits base classifiers each on random subsets of the original dataset and then aggregates their individual predictions to form a final prediction [50]. This approach helps reduce the variance of the model, as it combines multiple learners’ strengths and smooths out their predictions.
2.2.7 Support Vector Classifier (SVC)
Support Vector Classifier (SVC) is a powerful, flexible, and robust model for classification tasks [51]. It is based on the concept of decision planes that define decision boundaries. SVC aims to find a decision plane that maximizes the margin between different classes in the training data, making it particularly well-suited for binary classification problems.
@@ 2.3 FLOW-3D Simulation
FLOW-3D (v11.2) simulations are performed to accelerate process map development (Figure 1). FLOW-3D is a multiphysics simulation software produced by Flow Science, which provides more rapid estimations of the melt pool behavior than OpenFOAM. This reduction in computational expense is achieved by reducing the fluid flow within the vapor phase to estimates of the applied pressure and mass transfer at the surface of the melt pool. To create a dataset of FLOW-3D simulations, 283 SS316L single-track bare plate experiments are performed at varying processing parameters for a total length of 600 $\mu s$.
@@
$\nabla\cdot(\rho\vec{v})=0$ (7)
$\frac{\partial\vec{v}}{\partial t}+(\vec{v}\cdot\nabla)\vec{v}=-\frac{1}{\rho}\nabla\vec{P}+\mu\nabla^{2}\vec{v}+\vec{g}(1-\alpha(T-T_{m}))$ (8)
$\frac{\partial h}{\partial t}+\left(\vec{v}\cdot\nabla\right)h=\frac{1}{\rho}\left(\nabla\cdot k\nabla T\right)$ (9)
More specific information regarding the equations solved and physical phenomena considered during simulation can be found in [27, 52, 39]. This simulation is solved on a structured cartesian mesh, with mesh elements of size 10 $\mu m$. A complete list of the values for material parameters used to simulate the melt pool behavior can be found in the Appendix, in Table A.3.
@@ 2.4 Spatter Extraction and Analysis
Figure 2: A tracking algorithm is implemented to consistently identify spatter particles as they move through the domain. a) Before tracking, particles have no intrinsic ordering. After tracking, particles are identified by consistent indexes and can be correlated across timesteps. b) The tracking process is demonstrated for two sample particles $i$ and $j$, between the timesteps ${n}$ and ${n+1}$. During the tracking process, the velocity field of the particle is used to predict an estimated displacement $\delta x_{i}$ within $t_{n+1}-t_{n}=\Delta t$. Next, the actual particle positions at the next timestep, $i_{t_{n+1}}$, are correlated to the estimated displacement values $i^{\prime}_{t_{n+1}}$ based on a nearest neighbors assignment. Following this, each particle $i_{t_{n}}$ is linked to $i_{t_{n+1}}$, completing the tracking process.
@@
The OpenFOAM simulation data is processed to extract properties of both the ejected spatter and the melt pool. The cell-level simulation information is extracted for processing in Python, where each field is represented as a three-dimensional matrix defined by the mesh. As part of this information, three fields – $\alpha_{g}$, $\alpha_{s}$ and $\alpha_{l}$ denote the volume fraction of the gas phase, the solid phase, and the liquid phase respectively. To differentiate the powder bed, spatter, and melt pool from the gaseous phases of the simulation, the gas volume fraction $\alpha_{g}$ is truncated at 0.5, masking cells in the domain occupied by inert gas or metal vapor from the current analysis.
@@
Connected component analysis [53] is used to isolate individual spatter particles for further analysis. To do so, a binary field is constructed from $\alpha_{g}$ by applying a threshold at 0.5, assigning 1 to solid cells and liquid cells and 0 to gas phase cells. The spatter particles, powder bed and melt pool are then identified by applying the Union-Find algorithm to the binary array, which assigns unique labels to each disconnected object within the 3D domain. The largest contiguous object is identified as the composite melt pool and powder bed structure, and all other objects above a size-based noise threshold are considered to be spatter particles. Specifically, this threshold acts as a lower bound on the size of the particles, where particles comprised of less than eight connected cells are removed from the analysis. Following this processing step, the mesh elements that construct each particle can be used to extract particle-level field information.
@@
To extract the kinematics of the particle behavior, the positions of the particles must be tracked across the evolution of the simulation. During the connected components analysis, individual particles are identified without any intrinsic order. Therefore, an additional processing step is required to identify the mapping between particles at subsequent timesteps. We perform this tracking operation as a three-stage process – prediction, correlation and assignment. Given a set of $m$ particles at timestep $t_{n}$ and a set of $k$ particles at timestep $t_{n+1}$, the goal of this process is to associate each of the $m$ particles at $t_{n}$ with either the index of the corresponding particle at $t_{n+1}$, or a condition indicating that the particle trajectory has ended. The remaining particles at timestep $t_{n+1}$ that have not been associated with a particle at timestep $t_{n}$ are considered to have formed at timestep $t_{n+1}$.
@@
During the first phase of the tracking process, the velocity of each particle is used to estimate the displacement that occurs over a time $\Delta t$. Specifically, the average of the velocity fields for the cells comprising each particle is used to estimate the displacement $\delta\mathbf{x}_{i}$, assuming a constant velocity over the timestep $\Delta t$.
@@
To sample points from the melt pool surface, the volume fractions of the gas and liquid phases are used to define the melt pool morphology. Specifically, $\alpha_{g}$ is first used to threshold and identify the partially molten powder bed. From the points comprising the partially molten powder bed, the melt pool is identified by isolating values with a liquid volume fraction $\alpha_{l}$ greater than 0.5. The surface is then identified by finding the grid cells with the highest value of $\alpha_{g}$ within the points comprising the melt pool. During the sampling process, we extract the average of a specific field within a $n_{r}\times n_{r}\times n_{r}$ bounding cube centered at a specific point along the surface, only taking into account the points within the bounding cube that lie within the melt pool. We sample $n_{i}$ points from the melt pool at each timestep, where $n_{i}$ is the number of spatter particles ejected at time $i$. Each sample contains the mean of the position, velocity, density and temperature of the sampled region. Once this data has been sampled, the properties of the sampled regions are compared to the properties of the spatter particles.
Table 2: The melt pool dimensions comparison of experiment, FLOW-3D and OpenFOAM at a scan speed of 1 m/s. The experimental widths and depths listed are taken from cross-sections of the solidified melt track, and the reported error is two standard deviations
Power [W] 150 300 450
Experimental width $\mu m$ 105 ± 7.98 161 ± 9.00 181 ± 4.62
Flow3D width $\mu m$ 95 138 180
OpenFOAM width $\mu m$ 100 140 176
Experimental depth $\mu m$ 31.0 ± 3.32 78.9 ± 4.86 155 ± 8.58
Flow3D depth $\mu m$ 35 80 140
OpenFOAM depth $\mu m$ 38 73 152
@@ 3.1 Validation of Numerical Simulations
The simulation parameters and physics in the FLOW-3D and OpenFOAM models have been calibrated to match the experimental thermal data, following the methodology proposed by Myers and Quirarte et al. [52]. In these experiments, SS316L single-bead tracks were printed without powder using a TRUMPF TruPrint 3000 L-PBF machine. High-speed imaging of the melt pool was achieved using a Photron FASTCAM mini-AX200 color camera, capturing images at a rate of 22,500 frames per second, where each pixel corresponds to an area of 5.6 µm. These images were subsequently converted into pixel-by-pixel temperature data using the two-color method, which helps to reduce the impact of emissivity’s temperature dependence by using the intensity ratios from the camera’s red and green channels. Myers and Quirarte et al. [52] offer a detailed examination of the experimental methodologies employed. The dimensions of the melt pool are ascertained by sectioning, etching, and measuring the dimensions of the single-bead tracks.
@@
These comparisons are conducted at the laser powers that are within the range at which our OpenFOAM simulations for these studies were carried out (150 W, 300 W, and 450 W), with the scanning speed maintained at 1 m/s as shown in Table 1. At 100 W and 450 W while the scanning speed of 1 m/s, the width and depth of OpenFOAM simulation and experiment results are within the same range, while at 350 W and scanning speed of 1 m/s, there were slight variations. On comparing OpenFOAM meltpool dimensions to the lower bound of the width and depth of the experimental measurements at 300 W and 1 m/s, deviations of 7.8 % and and 1.3 % respectively were observed. These results show close agreements between experimental studies and OpenFOAM with respect to the measured depth and width. Although the deviation can be attributed to the laser in the experiment impacting the build plate at an angle of approximately 20 to accommodate the imaging apparatus [52]. Ultimately, developing a process window using Flow3D dataset via ML model trained on OpenFOAM dataset necessitates measurement agreements between both simulations. Slight variations with respect to the width and depth are observed with the highest being the measured depth at 450 W with 7.8 % decrease from OpenFOAM to Flow3D simulations.
@@
Further validation via comparison between the FLOW-3D, OpenFOAM data and experimental thermal profiles is displayed in Figure 3 for the sets of processing parameters, where the experimental measurements combine images taken at four different camera exposure times (1.05 $\mu$s, 1.99 $\mu$s, 6.67 $\mu$s, and 20 $\mu$s). Experimental temperature values below 2000 C are not visible via high-speed imaging could be due to the low intensity of the green and red thermal radiation at low temperatures. For the experimental temperature above 2500C, we were able to observe agreement among the experiment and the simulation for the larger section where the experimental dataset of the melt pool was collected.
@@
Figure 3: A comparison of the observed surface temperatures for single bead SS316L extracted via two-color thermal imaging as described in [52] to the equivalent surface temperatures extracted from FLOW-3D and OpenFOAM simulations. a) FLOW-3D, OpenFOAM and experimental measurements taken at P = 150 W, V = 1000 mm/s, a) FLOW-3D, OpenFOAM and experimental measurements taken at P = 300 W, V = 1000 mm/s. The error bars on the experimental data combine the uncertainty in emissivity and signal, as described in [52].
@@
Figure 4: Pairplot correlating spatial distribution of the meltpool and the spatter ejection through the histogram along the diagonal, Kernel Density Estimation (KDE) plots to the left of the diagonal and scatter plot to the right of the diagonal
@@
Figure 5: Pairplot correlating different measured properties of the meltpool and the spatter ejection through the histogram along the diagonal, Kernel Density Estimation (KDE) plots to the left of the diagonal and scatter plot to the right of the diagonal
@@
Figure 6: OpenFOAM simulation at 280 W and 1.5 m/s displaying spatter event at different timestep (a) spatter event 110 $\mu s$ (b) Spatter event at 440 $\mu s$
@@
Figure 7: Performance measurement of ML models on classifying spatter ejection and meltpool with the following features: x-axis, y-axis, z-axis, temperature, velocity, x-velocity, y-velocity, z-velocity, density and pressure (a) Accuracy measurement of the spatter and meltpool classification using different ML models on OpenFOAM dataset with all the features (b) Accuracy measurement of the spatter and the meltpool classification while position components are excluded; tuning the input features to match that can be extracted from FLOW-3D for spatter ejection prediction (c) Feature importance measurement of all the input features into ML model (d) Feature importance measurement of the input features excluding position components into ML model (e) Quantification of the correct and wrong predictions of spatter and meltpool when all features were used (f) Quantification of the correct and wrong predictions of spatter and meltpool when all features were used excluding position components

### v2

@@ 2.1 Simulation of LPBF via OpenFOAM
In equation 15, $m$ is the mass transfer between phases, $C^{\prime}$ is Lee’s model constant, $\alpha$ is the phase volume fraction, $T$ is the current temperature, $T_{A}$ is the solidification or melting temperature, $L$ is the latent heat. It should be noted that the same set of equations works for both solidification ($T$ $<$ $T_{A}$ and $C^{\prime}$ negative) and melting ($T$ $>$ $T_{A}$ and $C^{\prime}$ positive) [58].
Figure 2: Schematic illustration of the computational domain with random packing of stainless steel powder bed in for OpenFOAM simulation.
The dimension of the simulation domain is 1000 $\mu m$$\times$ 400 $\mu m$$\times$ 400 $\mu m$ as shown in Figure 2. A previous study by Liu et al. [59] indicated that to minimize or eliminate flowability uncertainty, it is preferable to reduce the powder size distribution to achieve near homogeneity. They found that the narrower the particle size distribution, the higher the mechanical resistance and hardness of the printed components. Another work by Young et al. [60] generated six sets of powder distribution that vary from 0 - 100 % of 15–25 $\mu m$ powder in a mixture with 35–48 $\mu m$ powder. They observed a general downward trend for spatter volume as the percentage of 15–25 $\mu m$ powder increased. Therefore, we choose to assume a uniform distribution of powder sizes for LPBF performance optimization. Powders of sizes 27 $\mu m$, a layer thickness of 81 $\mu m$, and a bare plate height of 142 $\mu m$ were used. Adaptive mesh was implemented in OpenFOAM where the mesh size along the x-axis and the y-axis is fixed at 2.5 $\mu m$ and 5 $\mu m$, respectively, while the z-axis varies from 2.5 $\mu m$ to 8.74 $\mu m$ to generate the total number of 2688000 cells.
Table 1: Material parameters implemented for the simulation of the SS316L melting process in OpenFOAM simulation [54, 61], while the subscript $s$ represents the solid phase, $l$ for the liquid phase, $v$ for the vapor and $n$ for the shielding gas nitrogen
@@ 2.2 FLOW-3D Simulation
FLOW-3D (v11.2) simulations are performed to accelerate the development of the process map (Figure 1). FLOW-3D is a multiphysics simulation software produced by Flow Science, which provides more rapid estimations of the melt pool behavior than OpenFOAM. This reduction in computational expense is achieved in part by abstracting the vapor phase dynamics to estimates of the applied pressure and mass transfer at the surface of the melt pool. To create a FLOW-3D simulation data set, 283 SS316L single-track bare plate experiments are performed at varying processing parameters for a total length of 600 $\mu s$.
@@
This simulation is carried out on a structured Cartesian mesh, with mesh elements sized at 10 $\mu m$. More specific information on the equations solved and the physical phenomena considered during the simulation can be found in [32, 62, 43, 46, 47].
@@ 2.3 Spatter Extraction and Analysis
Figure 3: A tracking algorithm is implemented to consistently identify spatter particles as they move through the domain. The tracking process is demonstrated for two sample particles $i$ and $j$, between the timesteps ${n}$ and ${n+1}$. During the tracking process, the velocity field of the particle is used to predict an estimated displacement $\delta x_{i}$ within $t_{n+1}-t_{n}=\Delta t$. Next, the actual particle positions at the next timestep, $i_{t_{n+1}}$, are correlated to the estimated displacement values $i^{\prime}_{t_{n+1}}$ based on a nearest neighbors assignment. Following this, each particle $i_{t_{n}}$ is linked to $i_{t_{n+1}}$, completing the tracking process.
@@
The OpenFOAM simulation data are processed to extract properties of both the ejected spatter and the melt pool. The cell-level simulation information is extracted for processing in Python, where each field is represented as a three-dimensional matrix defined by the mesh. As part of this information, three fields – $\alpha_{g}$, $\alpha_{s}$ and $\alpha_{l}$ denote the volume fraction of the gas phase, the solid phase, and the liquid phase, respectively. To differentiate the powder bed, spatter, and melt pool from the gaseous phases of the simulation, the gas volume fraction $\alpha_{g}$ is truncated at 0.5, masking cells in the domain occupied by inert gas or metal vapor from the current analysis.
@@
In this scenario, the ejected fluid spatter droplets as particles in our studies to denote a connected set of fluid cells that are disconnected from the larger melt pool. Connected component analysis [63] is used to isolate individual spatter particles for further analysis. To do so, a binary field is constructed from $\alpha_{g}$ by applying a threshold at 0.5, assigning 1 to solid cells and liquid cells and 0 to gas phase cells. The spatter particles, powder bed and melt pool are then identified by applying the Union-Find algorithm to the binary array, which assigns unique labels to each disconnected object within the 3D domain [64, 65]. The largest contiguous object is identified as the composite melt pool and powder bed structure, and all other objects above a size-based noise threshold are considered to be spatter particles. Specifically, this threshold acts as a lower bound on the size of the particles, where particles comprised of less than eight connected cells are removed from the analysis. Following this processing step, the mesh elements that construct each particle can be used to extract particle-level field information.
@@
To extract the kinematics of the particle behavior, the positions of the particles must be tracked across the evolution of the simulation. During the connected components analysis, individual particles are identified without any intrinsic order. Therefore, an additional processing step is required to identify the mapping between particles at subsequent timesteps. We perform this tracking operation as a three-stage process – prediction, correlation and assignment. Given a set of $m$ particles at timestep $t_{n}$ and a set of $k$ particles at timestep $t_{n+1}$, the goal of this process is to associate each of the $m$ particles at $t_{n}$ with either the index of the corresponding particle at $t_{n+1}$, or a condition indicating that the particle trajectory has ended. The remaining particles at timestep $t_{n+1}$ that have not been associated with a particle at timestep $t_{n}$ are considered to have formed at timestep $t_{n+1}$ (Figure 3).
@@
During the first phase of the tracking process, the velocity of each particle is used to estimate the displacement that occurs over a time $\Delta t$. Specifically, the average of the velocity fields for the cells comprising each particle is used to estimate the displacement $\delta\mathbf{x}_{i}$ (Equation 16), assuming a constant velocity over the timestep $\Delta t$. In Equation 16, $N_{s}$ refers to the number of cells comprising a particle, and $\mathbf{v_{s}}$ refers to the velocity of each individual cell.
@@
To sample points from the melt pool surface, the volume fractions of the gas and liquid phases are used to define the melt pool morphology. Specifically, $\alpha_{g}$ is first used to threshold and identify the partially molten powder bed. From the points comprising the partially molten powder bed, the melt pool is identified by isolating values with a liquid volume fraction $\alpha_{l}$ greater than 0.5. The surface is then identified by finding the grid cells with the highest value of $\alpha_{g}$ within the points comprising the melt pool. During the sampling process, we extract the average of a specific field within a $n_{r}\times n_{r}\times n_{r}$ bounding cube centered at a specific point along the surface, only taking into account the points within the bounding cube that lie within the melt pool, where $n_{r}$ is an arbitrarily selected parameter describing the size of the region investigated. We sample $n_{i}$ points from the melt pool at each timestep, where $n_{i}$ is the number of spatter particles ejected at time $i$. Each sample contains the mean of the position, velocity, density and temperature of the sampled region. Once this data has been sampled, the properties of the sampled regions are compared to the properties of the spatter particles.
2.4 Machine Learning Methodology
2.4.1 Data Collection and Hyper-Parameter Optimization
Our dataset was extracted from OpenFOAM simulation of two experiments, one conducted at a laser power of 545 W with a scan speed of 2 m/s, and another conducted at a laser power of 690 W and scan speed of 2 m/s. The spatter and the melt pool features were extracted using ParaView. The extracted features were used to validate our developed tracking algorithm, which is efficient in extracting not only spatter and melt pool properties but also the capability to track spatter trajectory. Detecting variations in the property distributions between the ejected spatter and the melt pool can enhance classification tasks. We implement different ML models as discussed in the Appendix A. to establish the performance trends for the spatter/melt pool classification task. From OpenFOAM experiments, we collected 488 data points, evenly split between the two classes (50 % spatter and 50 % melt pool). For the ML tasks, the dataset was divided into training (70 %) and testing (30 %) subsets using train test split, with a fixed random seed to ensure reproducibility of the results. Hyperparameter optimization was performed using GridSearchCV with 5-fold cross-validation, using the ROC-AUC score as the evaluation metric. The grid search explored a predefined hyperparameter space as listed on Table 2, the best parameters were identified by selecting the combination that maximized the average ROC-AUC score across the cross-validation folds. These parameters were then used to train the model in the training dataset. Predictions and probabilities were generated for both the training and testing sets. Model performance was evaluated using metrics including accuracy, F1-score, balanced accuracy, and ROC-AUC to assess the model’s predictive performance and generalization ability.
Table 2: Grid search hyperparameters range and the best value used for ML models.
Model Grid Search Parameters Used/Best Parameters
RF { ’n_estimators’: range(1, 100), ’max_depth’: range(3, 12), ’max_features’: [’sqrt’] } { ’criterion’: ’entropy’, ’max_depth’: 11, ’max_features’: ’sqrt’, ’n_estimators’: 90 }
GB { ’n_estimators’: range(50, 200, 50), ’learning_rate’: [0.01, 0.1, 0.2], ’max_depth’: range(3, 6), ’min_samples_split’: [2, 5, 10], ’min_samples_leaf’: [1, 3, 5] } { ’learning_rate’: 0.2, ’max_depth’: 4, ’min_samples_leaf’: 5, ’min_samples_split’: 2, ’n_estimators’: 150 }
Bagging BaggingClassifier( base_estimator=extra_tree, n_estimators=10, max_samples=0.8, max_features=0.8, bootstrap=True, random_state=42, n_jobs=-1)
ExtraTrees { ’n_estimators’: range(10, 100, 10), ’max_depth’: range(3, 10), ’max_features’: [’sqrt’, ’log2’, None], ’criterion’: [’gini’, ’entropy’] } { ’criterion’: ’entropy’, ’max_depth’: 9, ’max_features’: None, ’n_estimators’: 90 }
LGBM { ’n_estimators’: [100, 200], ’learning_rate’: [0.01, 0.1, 0.2], ’max_depth’: [-1, 10, 20], ’num_leaves’: [31, 50], ’min_child_samples’: [10, 20], ’subsample’: [0.8, 1.0], ’colsample_bytree’: [0.8, 1.0] } { ’colsample_bytree’: 1.0, ’learning_rate’: 0.2, ’max_depth’: -1, ’min_child_samples’: 20, ’n_estimators’: 200, ’num_leaves’: 31, ’subsample’: 0.8 }
KNN { ’n_neighbors’: range(1, 100), ’weights’: [’uniform’, ’distance’], ’p’: [1, 2] } { ’n_neighbors’: 16, ’p’: 1, ’weights’: ’distance’ }
SVC { ’svc__C’: [0.1, 1, 10, 100], ’svc__kernel’: [’linear’, ’rbf’, ’poly’], ’svc__gamma’: [’scale’, ’auto’], ’svc__degree’: [2, 3] } { ’svc__C’: 10, ’svc__degree’: 2, ’svc__gamma’: ’scale’, ’svc__kernel’: ’rbf’ }
2.4.2 SHapley Additive exPlanations as an Explainable AI
SHapley Additive exPlanations (SHAP) implement a game-theoretic approach to describe how the ML models make predictions. It is based on the concept of Shapley values from cooperative game theory, where contributions of individual players (features) to the overall game outcome (model prediction) are calculated fairly [66]. SHAP values assign each feature a contribution value for every prediction, ensuring both global and local interpretability. The method guarantees local accuracy (the sum of SHAP values equals the model prediction) and consistency (a feature with increased contribution in a newer model will not have a lower SHAP value).
Mathematically, the SHAP value for a feature $i$ is given by:
$\phi_{i}=\sum_{S\subseteq N\setminus\{i\}}\frac{|S|!(|N|-|S|-1)!}{|N|!}\big{[}f(S\cup\{i\})-f(S)\big{]},$ (18)
where $S$ is a subset of features, $N$ is the set of all features, and $f(S)$ represents the model prediction with features in $S$. This ensures that SHAP fairly distributes the prediction among features by evaluating all possible subsets.
SHAP provides two key benefits: (1) Global interpretability, where aggregated SHAP values reveal feature importance across the dataset, and (2) Local interpretability, where SHAP values explain individual predictions by quantifying each feature’s contribution.
Partial Dependence Plots (PDPs) provide a visual representation of how one or two features influence model predictions, independent of other features. For a single feature $X_{j}$, the partial dependence function is defined as [67]:
$\text{PDP}(X_{j})=\frac{1}{n}\sum_{i=1}^{n}f(X_{i,-j},X_{j})$ (19)
where $X_{i,-j}$ denotes all features except $X_{j}$. Single-feature PDPs illustrate the marginal effect of a feature, while pairwise PDPs show the joint effect of two features on predictions. PDPs are particularly useful for validating model interpretability and identifying nonlinear relationships.
Figure 4: Mesh dependence studies on the temperature profile (a) FLOW-3D (b) OpenFOAM
@@ 3.1 Validation of Numerical Simulations
Mesh independence experiments were carried out on OpenFOAM and FLOW-3D as shown on Figure 4. Mesh sizes of 5 $\mu m$ and 10 $\mu m$ were compared in FLOW-3D, while mesh sizes of 2.5 $\times$ 5 $\times$ 2.5-8.74 $\mu m$, 5 $\mu m$ and 10 $\mu m$ were compared in OpenFOAM. Close matches were observed in OpenFOAM and FLOW-3D, indicating that the mesh resolves heat transfer at the chosen resolution across the mesh sizes investigated.
@@
The simulation parameters and physics in the FLOW-3D and OpenFOAM models have been calibrated to match the experimental thermal results, following the methodology proposed by Myers and Quirarte et al. [62]. In these experiments, SS316L single-bead tracks were printed without powder using a TRUMPF TruPrint 3000 L-PBF machine. High-speed imaging of the melt pool was achieved using a Photron FASTCAM mini-AX200 color camera, capturing images at a rate of 22,500 frames per second, where each pixel corresponds to an area of 5.6 $\mu m$. These images were subsequently converted into pixel-by-pixel temperature data using the two-color method, which helps to reduce the impact of emissivity’s temperature dependence by using the intensity ratios from the camera’s red and green channels. Myers and Quirarte et al. [62] provide detailed experimental methodologies and results. The dimensions of the melt pool are determined by sectioning, etching, and measuring the dimensions of the single-bead tracks.
@@
These comparisons are conducted at the laser powers that are within the range at which our OpenFOAM simulations for these studies were carried out (150 W, 300 W, and 450 W), with the scanning speed maintained at 1 m/s as shown in Table 1. At 100 W and 450 W while the scanning speed is 1 m/s, the width and depth of OpenFOAM simulation and experiment results are within the same range, while at 350 W and scanning speed of 1 m/s, there were slight variations. On comparing OpenFOAM melt pool dimensions to the lower bound of the width and depth of the experimental measurements at 300 W and 1 m/s, deviations of 7.8 % and and 1.3 % respectively were observed. These results show close agreements between experimental studies and OpenFOAM with respect to the measured depth and width. The deviation may be due to the experimental settings of the laser orientation as it was positioned at an angle of approximately 20 during the interaction with the build plate for the accommodation of the imaging apparatus [62]. Ultimately, developing a process window using the Flow3D dataset via the ML model trained on the OpenFOAM dataset requires measurement agreements between both simulations. Slight variations with respect to the width and depth are observed with the highest being the measured depth at 450 W with 7.8 % decrease from OpenFOAM to Flow3D simulations.
Table 3: The melt pool dimensions comparison of experiment, FLOW-3D and OpenFOAM at a fixed speed of 1 m/s.The experimental widths and depths provided are generated from the solidified melt track, while the reported errors are in two standard deviations.
Power [W] 150 300 450
Experimental width ($\mu m$) 105 ± 7.98 161 ± 9.00 181 ± 4.62
Flow3D width ($\mu m$) 95 138 180
OpenFOAM width ($\mu m$) 100 140 176
Experimental depth ($\mu m$) 31.0 ± 3.32 78.9 ± 4.86 155 ± 8.58
Flow3D depth ($\mu m$) 35 80 140
OpenFOAM depth ($\mu m$) 38 73 152
@@
Further validation via comparison between the FLOW-3D, OpenFOAM data and experimental temperature profiles is displayed in Figure 3 for the set of process conditions, where the experimental measurements combine images captured at different camera exposure times (1.05 $\mu$s, 1.99 $\mu$s, 6.67 $\mu$s, and 20 $\mu$s). Experimental temperature values below 2000 C are not visible via high-speed imaging could be due to the low intensity of the green and red thermal radiation at low temperatures. The deviation in the experimental measurements is due to the plume emission and transmission above the melt pool that interferes with temperature measurement and obscure the front and side melt pool boundaries with temperature. However, a close range and similar rise-and-fall profile can be observed between the experimental and simulation results, particularly for the larger section where the experimental melt pool dataset was collected [62]. In addition to validating the melt pool dimensions and the temperature distribution, the dynamics of the spatter simulated in OpenFOAM were validated using the experimentally reported velocities by Ly et al [19]. The velocity distribution from OpenFOAM was extracted (Figure 6a) and compared to the values reported by Ly et al. on the y-axis of Figure 6b, which ranged from 2.5 m/s to 20 m/s for recoil pressure spatter and hot spatter as shown Figure 6b, demonstrating agreement with the OpenFOAM velocity distribution.
@@
Figure 5: Implementation of the two-color thermal imaging method as described in [62] for the comparison of the observed surface temperatures for SS316L with the corresponding surface temperatures obtained from modelling packages: a) Measurements taken at P = 150 W, V = 1000 mm/s b)Measurements taken at P = 300 W, V = 1000 mm/s. The error bars from the experimental measurements represent combined uncertainties in emissivity and signal, as detailed in [62].
Figure 6: Spatter dynamic validation via the magnitude of the velocity (a) OpenFOAM velocity distribution (b) Reported velocity by Ly et al. [19].
@@
Figure 7: Pairplot correlating spatial distribution of the melt pool and the spatter ejection through the histogram along the diagonal, Kernel Density Estimation (KDE) plots to the left of the diagonal and scatter plot to the right of the diagonal .
@@
Figure 8: Pairplot correlating different measured properties of the melt pool and the spatter ejection through the histogram along the diagonal, Kernel Density Estimation (KDE) plots to the left of the diagonal and scatter plot to the right of the diagonal.
@@
Figure 9: OpenFOAM simulation at 509 W and 1.5 m/s displaying spatter event at different timestep (a) spatter event 110 $\mu s$ (b) Spatter event at 440 $\mu s$.
@@
Figure 10: Performance measurement of ML models on classifying spatter ejection and melt pool with the following features: x-axis, y-axis, z-axis, temperature, velocity, x-velocity, y-velocity, z-velocity, density and pressure (a) ROC_AUC score (b) Balanced_Accuracy score (c) F1 score (d) Accuracy (e) Feature importance measurement from RF model (f) Quantification of the correct and wrong predictions of spatter and melt pool from RF model.
Figure 11: Performance measurement of ML models on classifying spatter ejection and melt pool with the following features: temperature, velocity, x-velocity, y-velocity, z-velocity, density and pressure (a) ROC_AUC score (b) Balanced_Accuracy score (c) F1 score(d) Accuracy (e) Feature importance measurement from RF model (f) Quantification of the correct and wrong predictions of spatter and melt pool from RF model.
Figure 12: SHAP analysis on RF model in classifying spatter ejection and melt pool with the following features: temperature, velocity, x-velocity, y-velocity, z-velocity, density and pressure (a) SHAP feature important (b) The SHAP summary plot with a color bar (c) pressure PDP (d) z-velocity PDP (e) temperature PDP (f) x-velocity PDP (g) Velocity magnitude PDP (h) Density PDP (i) y-velocity PDP
Figure 13: SHAP analysis on RF model in classifying spatter ejection and melt pool with the following features: temperature, velocity, x-velocity, y-velocity, z-velocity, density and pressure (a) SHAP feature important (b) Balanced_Accuracy score (c) z-axis PDP (d) pressure PDP (e)Partial dependence plot on y-axis PDP (f) z-velocity PDP (g) Velocity magnitude PDP(h) Temperature PDP (i)x-velocity PDP (j) x-axis PDP (k)Density PDP(l) y-velocity PDP